A 2.007 gram sample of a hydrocarbon is combusted to give 1.389 grams of H[sub]2[/sub]O and 6.785 grams of CO[sub]2[/sub]. What is the empirical formula of the compund?
1
Expert's answer
2011-05-11T04:46:26-0400
CnHm + xO2 = nCO2 + m/2H2O

2.007 g 6.785 g 1.389 g


n(H2O) = 1.389/18 = 0.0771mole
n(CO2) = 6.785/44 = 0.1542mole
n/m = n(CO2)/(2*n(H2O)) = 2 therefore it is alkene CnH2n
We cannot say more, because carbon- hydrogenatoms ratio is constant.